OpenStudy (anonymous):

Find the surface area of a cylinder with a radius of 12.2mm and a height of 5.9mm to the nearest square unit. A.1,375.1 mm2 B.1,386.7 mm2 C.1,389.3 mm2 D.1,408.5 mm2

OpenStudy (anonymous):

Is it closed up and down?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

\[ Area = 2 \pi r h + 2 \pi r^2=1386.75 \] if \[\pi \approx 3.14\]

OpenStudy (anonymous):

Ohhh , okayy thankyou(:

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

No im good lol, thanks tho (: